A nylon wire with a cross-sectional area of 8.0 x10-7 m2 has a Young modulus of 3.5 x108Pa. A force of 100N is applied to the wire Calculate the stress on the wire
1
Expert's answer
2020-11-17T12:58:08-0500

We can find the stress on the wire from the formula:


"\sigma=\dfrac{F}{A},"

here, "\sigma" is the stress on the wire, "F=100\ N" is the force applied to the wire, "A=8.0\cdot10^{-7}\ m^2" is the cross-sectional area of the wire.

Then, from this formula we can calculate the stress on the wire:


"\sigma=\dfrac{100\ N}{8.0\cdot10^{-7}\ m^2} =125\cdot10^6\ Pa =125\ MPa."

Answer:

"\sigma=125\cdot10^6\ Pa=125\ MPa."